OpenStudy (anonymous):

What is the exact value of 3q^8 over sqrt q^19

OpenStudy (anonymous):

\[\frac{3Q^{8}}{\sqrt{Q^{19}}}\] Simplify the denominator.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Rewrite it using these two general formulas: \[\sqrt{AB} = \sqrt{A} \sqrt{B}\] \[A^{X+Y} = A^{X}*A^{Y}\]

OpenStudy (anonymous):

Which is ??? I'm really bad at this..

OpenStudy (anonymous):

\[A^{T} = A^{X+Y}\]

OpenStudy (anonymous):

whats substituted into the equation?

OpenStudy (anonymous):

\[Q^{19} = Q^{X+Y} = Q^{X}*Q^{Y}\] \[\sqrt{Q^{19}} = \sqrt{Q^{X+Y}} = \sqrt{Q^{X}*Q^{Y}} = \sqrt{Q^{X}}*\sqrt{Q^{Y}}\]

OpenStudy (anonymous):

Of course you need to figure out X and Y. You can always reduce it bit by bit, until you can't reduce it anymore.

OpenStudy (anonymous):

Example: \[\sqrt{Q^{5}} = \sqrt{Q^{2+3}} = \sqrt{Q^{2}*Q^{3}} = \sqrt{Q^{2}}*\sqrt{Q^{3}} = Q*\sqrt{Q^{3}}\] \[Q*\sqrt{Q^{3}} = Q*\sqrt{Q^{2+1}} = Q*\sqrt{Q^{2}*Q^{1}} = Q*\sqrt{Q^{2}}*\sqrt{Q^{1}} = Q*Q*\sqrt{Q^{1}} = Q^{2}\sqrt{Q}\] The same thing can be done quicker as: \[\sqrt{Q^{5}} = \sqrt{Q^{4+1}} = \sqrt{Q^{4}*Q^{1}} = \sqrt{Q^{4}}*\sqrt{Q^{1}} = Q^{2}*\sqrt{Q^{1}} = Q^{2}\sqrt{Q}\]

OpenStudy (anonymous):

so the answer is ?? lol

OpenStudy (anonymous):

Do you just want the answer, or do you want to know how to solve it for yourself?

OpenStudy (anonymous):

The answer please, it's a final test.